At the start of year 2015, a certain city has a population of 15,000. If the population increases by 2% every year, in how many years will the population be more than triple the original number?


1
Expert's answer
2021-09-10T00:28:28-0400

Solution:

"P=P_0(1.02)^t" ...(i)

Here, P = Population after t years

And "P_0" = Initial population.

Given that "P_0=15000,P=3\times15000=45000"

We need to find t.

Putting all these values in (i).

"45000=15000(1.02)^t
\\\Rightarrow3=(1.02)^t
\\\Rightarrow \log 3=t \log1.02
\\\Rightarrow t=55.47"

Thus, after 56 years, the population be more than triple the original number.